OpenStudy (timaashorty):

Can anyone take me through a problem about percentiles step by step? Please? (:

OpenStudy (doc.brown):

I love percent problems 0.5 out of 1

OpenStudy (doc.brown):

How much is that out of 100?

OpenStudy (timaashorty):

lol 50%

OpenStudy (timaashorty):

Will you help me? c;

OpenStudy (doc.brown):

Show your work:\[\frac{0.5}{1}\times\frac{100}{100}=\frac{50}{100}\]

OpenStudy (doc.brown):

What is your concern?

OpenStudy (timaashorty):

I didn't had to work it out. it's obvous that 0.5 is 1/2 of 1. which means 50% ;P

OpenStudy (timaashorty):

Find the values of the 30th and 90th percentiles of the data. 129, 113, 200, 100, 105, 132, 100, 176, 146, 152

OpenStudy (timaashorty):

That^ Do I need to put them in order first or what?

OpenStudy (doc.brown):

Statistics?

OpenStudy (timaashorty):

yeah

OpenStudy (timaashorty):

Do you know it?

OpenStudy (doc.brown):

Ok yes, put them in order. the 30th percentile is the top 70%

OpenStudy (doc.brown):

You have ten numbers, I counted, which means you want the third biggest to the biggest then you want the 9th biggest and the biggest, for the top 10% (90% and higher)

OpenStudy (timaashorty):

\[100 \] \[100 \] \[105 \] \[113\] \[ 129 \] \[132 \] \[ 146 \] \[152 \] \[ 176 \] \[200 \]

OpenStudy (timaashorty):

wait so 3rd biggest is 105 and 9th biggest is 176

OpenStudy (timaashorty):

I have that as one of my answer choices. is that all we do ?

OpenStudy (doc.brown):

Yeah, let me put it this way: If a number is in the 90th percentile, it is higher than 90% of the other scores

OpenStudy (timaashorty):

and if it's the 3oth percentile it's 30 % of the other numbers?

OpenStudy (doc.brown):

Higher than 30%

OpenStudy (doc.brown):

If you are in the 51st percentile, you are doing better than half your class.

OpenStudy (timaashorty):

Thank very much! You really helped [: